# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v1

@@ VIII-A Roadmap Construction
We present the roadmap construction method in Algorithm 4. The set of contact configurations $R$ in Algorithm 4 are generated by solving (40) for $q^{\mathrm{o}}$’s from a set of stable object configurations which sufficiently cover the object workspace. For each pair of configurations $(q_{i},q_{j})$ in $R$, we first try to reach $q^{\mathrm{o}}_{j}$ from $q_{i}$ by running MPC (Algorithm 4), and then reach $q^{\mathrm{a}}_{j}$ by standard collision-free motion planning (Algorithm 4). This procedure is illustrated in Figure 19. We repeat this procedure and add all successful connections to the roadmap.
@@ 3 $V\leftarrow\emptyset$, $E\leftarrow\emptyset$;
4 Generate contact configurations $R\coloneqq\{(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i})\}_{i=1}^{m}$ ;
5 for ordered pair $q_{i}\coloneqq(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i}),q_{j}\coloneqq(q^{\mathrm{o}}_{j},q^{\mathrm{a}}_{j})$ in $R$ do
@@ 7       $q_{H:M},u_{H:M-1}\leftarrow\mathbf{CollisionFree}(q_{H},q_{j})$;
Consider the cube configuration in Figure 20a, which corresponds to the identity rotation (i.e. no rotation). By leveraging symmetry, we can reach its 24 rotational symmetries using just three basic operations: i) yaw from $0^{\circ}$ to $90^{\circ}$, ii) yaw from $0^{\circ}$ to $-90^{\circ}$ and iii) pitch from $0^{\circ}$ to $90^{\circ}$. These three operations require just five grasps in total: 2 for each yaw and 1 for the pitch (see Figure 18). The grasps take about 5 minutes to generate (Table VI). Moreover, connecting each pair of grasping configurations with MPC and collision-free motion planning takes a few seconds (Table III), keeping the total roadmap construction time under 10 minutes. Parallelizing these steps can further reduce computation time.
To verify the robustness of the constructed roadmap, we conducted a random walk on the roadmap on hardware, and recorded 150 successful consecutive edge transitions before hardware failure occurred (the hand overheated). Recording of this experiment can be found in the supplementary video.
Figure 20: A complete path generated by our roadmap-based global planner. The object configuration in each frame is highlighted at the lower-left corner. (a) shows the starting configuration, which in this example is a vertex already in the roadmap. (b) can be reached from (a) with a $-90^{\circ}$ yaw. (c) has the same object configuration as (b), but the hand has repositioned for a $90^{\circ}$ pitch. (d) is system configuration after the pitch. (e) shows the system reaching the goal configuration from (d). The path (a)-(d) is part of the roadmap and generated offline. (d)-(e) is generated online using collision-free planning and local MPC.
VIII-B Inference on a Roadmap
@@
IX Conclusion
@@
Have we solved the problem of planning and control through contact dynamics? Locally—on CQDC dynamics—the proposed MPC with Contact Trust Region constraints can achieve small tracking errors for the vast majority of the goals we sampled based on local reachability criteria. However, we do not understand this problem nearly as well as we do the simple pendulum: not only do we lack solid explanations for the small but non-zero number of planner failures, but we also have yet to fully understand the different roles played by feasibility constraints on IiwaBimanual vs AllegroHand.
@@
Much more remains to be understood for second-order dynamics, both in simulation and on hardware. In particular, keeping the robot in contact with the object, without exploiting the artifact of CQDC dynamics, remains one of the biggest unsolved challenges. Incorporating the CTR constraints in MPC greatly alleviates the problem of lost contact, and the initial guess heuristics empirically helps a bit more. However, we occasionally still observe loss of contact, and its root cause remains unclear.
@@
Nevertheless, the tools presented in this paper already enable capabilities that were previously out of reach for model-based methods. By accounting for the unilateral nature of contact, our contact trust region makes it possible to apply a broad range of robotics algorithms to contact-rich manipulation problems. We hope the MPC, grasp synthesis, and roadmap-based global planning methods introduced here are only a small sample of the many contact-rich manipulation algorithms yet to come.
X Appendix
X-A Derivation of (15)
@@
We note that the matrix equation (15) is not invertible unless the primal and dual cones defined in (5) were one-dimensional, e.g. $\mathcal{K}=\mathbb{R}_{+}$ and $\mathcal{K^{\star}}=\mathbb{R}_{+}$. This holds in the quadratic programming formulation of contact dynamics [10] where constraints are simple inequalities instead of conic inequalities. For higher-dimensional conic constraints such as (5), differentiating (13) and (12) (instead of (9)) together would produce an invertible systems of equation to which the implicit function theorem can be applied. However, since deriving the derivatives involves complicated matrix differential calculus [73], we omit its derivation in this paper.
X-B Proof of Lemma 2
@@
X-C Goal Generation for Local MPC
To generate goals that are considered more locally reachable yet far enough to be challenging for the planner, we utilize the following schemes. For the IiwaBimanual environment, we first prescribe some initial state $\bar{q}$ by sampling contact points on the surface of the object and solving inverse kinematics; then, the goals are sampled from the boundary of the action-only object motion set $\mathcal{M}^{\mathcal{A},\mathrm{o}}_{\mathbf{\Sigma},\kappa}(\bar{q},\bar{u})$ with $\bar{u}=\bar{q}^{\mathrm{a}}$. To make the goals sufficiently challenging, we use large ellipsoid radius for the purpose of sampling goals. The resulting goals have maximum $120^{\circ}$ rotation and $0.4m$ translation from the initial state (see Figure 9 for the distrubtion of goals). Similarly, we prescribe some grasp for AllegroHand and sample orientations that have about a $60^{\circ}$ difference from the initial state according to the axis-angle metric. With the above procedure, we sampled $1233$ pairs of initial states and goals for IiwaBimanual and $1000$ pairs for AllegroHand.
We summarize in Table VII the average amount of translation and rotation between the initial and goal object configurations generated here. These would be the object pose tracking errors if the planning algorithm being evaluated did nothing.
@@ X-C Goal Generation for Local MPC
Mean $\|p_{\text{goal}}-p_{\text{0}}\|$ [mm] 152 16
Mean $|\Delta\theta(Q_{\text{goal}},Q_{\text{0}})|$ [mrad] 356 788
TABLE VII: Mean translation and rotation differences for pairs of initial and goal object states generated in this section.
X-D Effect of Planning Horizon for Local MPC
We study how the planning horizon $T$ used in (36) affects the MPC errors. Specifically, for the IiwaBimanual system and the corresponding goals generated using the scheme in Section X-C, we run Algorithm 2 with planning horizon from $1$ to $5$, a maximum iteration of $2$ ($T=1\dots 5$, $n_{\mathrm{max}}=2$ in Algorithm 1), and a MPC rollout horizon $H$ of $10$.
The results are summarized in Table VIII. Surprisingly, we found that the planning horizon $T$ has little to no effect on the final error for MPC: our method is able to reach the goals within a tight tolerance for all values of $T$. We believe that this is due to two reasons: i) we have chosen the initial configurations and goals to be reasonably reachable using our trust region, and ii) the CQDC model allows large change in configurations within a single step, so we are able to get away with significantly less horizon compared to other MPC tasks such as dynamic locomotion.
@@ X-D Effect of Planning Horizon for Local MPC
TABLE VIII: Translation and rotation Errors for IiwaBimanual under different planning horizon $T$. The first row is the mean; the second row (numbers in brackets) is the standard deviation.
X-E Hyperparameters for 2nd-order Dynamics Stabilization
We use the same $\mathbf{CtrTrajOpt}$ parameters from Table I. Other important MPC hyperparameters are summarized in Table IX. For closed-loop IiwaBimanual experiments, we use larger $N$ for goals that are further away, while keeping $N\times H$ around $25$. In contrast, as slippage happens a lot more often on the AllegroHand system, we keep $N$ constant for all goals to give MPC more chances to recover.
$N$ (Number of re-plans) $H$ (MPC rollout horizon)

### v3

@@ 8.1 Roadmap Construction
We present the roadmap construction method in Algorithm 4. The set of grasping configurations $\mathcal{G}$ in Algorithm 4 are generated by solving (40) for $q^{\mathrm{o}}$’s from a set of stable object configurations which sufficiently cover the object workspace. For each pair of configurations $(q_{i},q_{j})$ in $\mathcal{G}$, we first try to reach $q^{\mathrm{o}}_{j}$ from $q_{i}$ by running MPC (Algorithm 4), and then reach $q^{\mathrm{a}}_{j}$ by standard collision-free motion planning (Algorithm 4). This procedure is illustrated in Figure 20. We repeat this procedure and add all successful connections to the roadmap.
@@ 3 $V\leftarrow\emptyset$, $E\leftarrow\emptyset$;
4 Generate contact configurations $\mathcal{G}\coloneqq\{(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i})\}_{i=1}^{m}$ ;
5 for ordered pair $q_{i}\coloneqq(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i}),q_{j}\coloneqq(q^{\mathrm{o}}_{j},q^{\mathrm{a}}_{j})$ in $\mathcal{G}$ do
@@ 7 $q_{H:M},u_{H:M-1}\leftarrow\mathbf{CollisionFree}(q_{H},q_{j})$;
Consider the cube configuration in Figure 21a, which corresponds to the identity rotation (i.e. no rotation). By leveraging symmetry, we can reach its 24 rotational symmetries using just three basic operations: (i) yaw from $0^{\circ}$ to $90^{\circ}$, (ii) yaw from $0^{\circ}$ to $90^{\circ}$ and (iii) pitch from $0^{\circ}$ to $90^{\circ}$. These three operations require just five grasps in total: 2 for each yaw and 1 for the pitch (see Figure 19). The grasps take about 5 minutes to generate (Table 8). Moreover, connecting each pair of grasping configurations with MPC and collision-free motion planning takes a few seconds (Table 5), keeping the total roadmap construction time under 10 minutes. Parallelizing these steps can further reduce computation time.
Figure 21: A complete path generated by our roadmap-based global planner. The object configuration in each frame is highlighted at the lower-left corner. (a) shows the starting configuration, which in this example is a vertex already in the roadmap. (b) can be reached from (a) with a $90^{\circ}$ yaw. (c) has the same object configuration as (b), but the hand has repositioned for a $90^{\circ}$ pitch. (d) is system configuration after the pitch. (e) shows the system reaching the goal configuration from (d). The path (a)-(d) is part of the roadmap and generated offline. (d)-(e) is generated online using collision-free planning and local MPC.
To verify the robustness of the constructed roadmap, we conducted a random walk on the roadmap on hardware, and recorded 150 successful consecutive edge transitions before hardware failure occurred (the hand overheated). Recording of this experiment can be found in the supplementary video. The high robustness of the roadmap can be attributed to (i) the simplicity of the cube’s geometry, whose symmetry makes connecting nodes with MPC a lot easier; (ii) the caging property of the grasps (i.e. nodes), which consistently guides the cube back to the center of the palm.
8.2 Inference on a Roadmap
@@
8.3 Limitations
The primary limitation of our global planner is that its success depends on a connection recipe that is not yet general-purpose. Our current implementation heavily leverages the specific kinematics of the AllegroHand and the symmetries of the cube.
Our method for connecting two roadmap nodes, $(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i})\in\mathcal{G}$ and $(q^{\mathrm{o}}_{j},q^{\mathrm{a}}_{j})\in\mathcal{G}$, involves two steps: (i) an MPC plan moves the object from $q^{\mathrm{o}}_{i}$ to $q^{\mathrm{o}}_{j}$, and (ii) a collision-free motion plan repositions the hand from the hand configuration reached by MPC to $q^{\mathrm{a}}_{j}$. Both steps are made easy by carefully selecting nodes that correspond to the 24 stable, symmetric rotations of the cube resting flat on the palm.
Our choice of nodes trivializes the first step, as connecting any $q^{\mathrm{o}}_{i}$ to $q^{\mathrm{o}}_{j}$ only needs to roll, pitch or yaw the cube by 90 degrees. However, for generic object shapes without exploitable symmetry, connectivity of $q^{\mathrm{o}}_{j}$ from $(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i})$ needs to be computed more explicitly. One possible strategy is to build a convex inner approximation of the set of object poses reachable from $(q^{\mathrm{o}}_{i},q^{\mathrm{a}}_{i})$ using a method similar to Petersen and Tedrake (2023), and attempt to connect only if $q^{\mathrm{o}}_{j}$ belongs to the approximated reachable set.
The second step critically assumes that the object remains stable and stationary at $q^{\mathrm{o}}_{j}$ while the hand moves. This assumption, however, would likely fail for objects without nice symmetries or for tasks requiring grasps on unstable object poses, as the collision-free planner could no longer assume a static object. A more general method would need to explicitly reason about maintaining object stability during these hand-repositioning phases, such as a finger-gaiting strategy where a subset of fingers hold the object still while the other fingers reposition.
9 Discussion and Conclusion
9.1 Synergies with Learning-based Methods
Our contact-rich planners have strong connections to modern robot learning, particularly in addressing key bottlenecks in (i) large-scale imitation learning (IL) and (ii) reinforcement learning (RL).
9.1.1 Data Synthesis and Augmentation for IL
Recent advances in large-scale imitation learning are impressive (Black et al., 2024; TRI et al., 2025), but their performance hinges on large-scale, embodiment-specific datasets. This data is typically collected via human teleoperation, which presents a significant bottleneck for two reasons: not only is teleoperation time-consuming and expensive, teleoperation interfaces are also often limited to controlling robot end effectors, making it challenging for humans to provide high-quality demonstrations for tasks involving whole-body or in-hand contact.
Our global planner directly addresses this challenge by serving as an automated expert data generator. By leveraging a model-based planner, we can create vast datasets of successful trajectories for complex, contact-rich tasks that are hard to teleoperate. For instance, we can generate synthetic expert demonstrations for tasks like IiwaBimanual and AllegroHand (Zhu et al., 2025) to train a goal-conditioned diffusion policy (Chi et al., 2024). The resulting policy could achieve a high success rate with zero human demonstrations and would be more computationally efficient at runtime than using the planner, as it wouldn’t need to replan from scratch.
Beyond replacing human data, our planners can also be used to augment it. Human demonstrations often cover a narrow distribution of successful trajectories. A policy trained solely on this data can be brittle and fail when faced with small perturbations. Our local MPC planner can significantly enhance policy robustness. Following the approach in Yang et al. (2025), we can take a human demonstration, inject noise at various points, and use a local MPC planner to generate a corrective action that brings the robot back toward the goal. This process creates a rich dataset of near-failure states and successful recoveries, teaching the learned policy how to handle perturbations far more effectively than a policy trained only on the original, unperturbed data.
9.1.2 Efficient Exploration for RL
Our global planner can also serve as efficient exploration mechanisms for RL. Despite many algorithmic advances, efficient exploration still remains a significant challenge for RL. For problems in Figure 16, for example, a naive Gaussian sampling exploration scheme would have a really hard time finding an effective policy. In order to bypass this difficulty, many practical RL algorithms leverage expert demonstrations either as rewards (Rajeswaran et al., 2018; Peng et al., 2018) or as random resets during training (Song et al., 2023) to guide RL exploration towards relevant states that are more beneficial to accomplish the task.
Both (Khandate et al., 2024) and (Barreiros et al., 2025) have successfully explored the effectiveness of using model-based planners such as (Pang et al., 2023) to vastly improve the sample efficiency of exploration in RL. We hypothesize that the stable states in our roadmap, for example, can serve as an extremely efficient exploration distribution in RL.
9.2 Concluding Remarks and Limitations
@@
Have we solved the problem of planning and control through contact dynamics? Locally—on CQDC dynamics—the proposed MPC with Contact Trust Region constraints can achieve small tracking errors for the vast majority of the goals we sampled based on local reachability criteria. The local MPC can be readily integrated with a sampling-based planner to enable global search, although more work is necessary to generalize our global planner beyond the cube.
We demonstrated our planner’s efficacy on two distinct platforms: IiwaBimanual and AllegroHand. These examples were intentionally chosen to highlight the method’s generalizability, as they represent significantly different scales, kinematic structures, and contact scenarios—from whole-body manipulation to dexterous, in-hand motion. Since our method only needs robot and object URDFs, it can be readily extended to other contact-rich systems, including different hands, grippers, or even legged robots (for which we may need to add balancing constraint to trust regions).
However, we do not understand this problem nearly as well as we do the simple pendulum: not only do we lack solid explanations for the small but non-zero number of planner failures, but we also have yet to fully understand the different roles played by feasibility constraints on IiwaBimanual vs AllegroHand.
@@
Moreover, much more remains to be understood for second-order dynamics, both in simulation and on hardware. In particular, keeping the robot in contact with the object, without exploiting the artifact of CQDC dynamics, remains one of the biggest unsolved challenges. Incorporating the CTR constraints in MPC greatly alleviates the problem of lost contact, and the initial guess heuristics empirically helps a bit more. However, we occasionally still observe loss of contact, and its root cause remains unclear.
@@
Nevertheless, the tools presented in this paper already enable capabilities that were previously out of reach for model-based methods. By accounting for the unilateral nature of contact, our contact trust region makes it possible to apply a broad range of robotics algorithms to contact-rich manipulation problems. We hope the MPC, grasp synthesis, and roadmap-based global planning methods introduced here are only a small sample of the many contact-rich manipulation algorithms yet to come. {acks} We thank Stephen Proulx, Xinpei Ni, and Velin Dimitrov for building the cube with active OptiTrack markers, and Giovanni Remigi for making the project website.
This work is funded by the Defense Science & Technology Agency Award DST00OECI20300823, Amazon Award PO#2D-15694085, and the Boston Dynamics AI Institute Award Agmd Dtd 8/1/2023.
Appendix A Appendix
A.1 Derivation of (15)
@@
We note that the matrix equation (15) is not invertible unless the primal and dual cones defined in (5) were one-dimensional, e.g. $\mathcal{K}=\mathbb{R}_{+}$ and $\mathcal{K^{\star}}=\mathbb{R}_{+}$. This holds in the quadratic programming formulation of contact dynamics (Pang and Tedrake, 2021) where constraints are simple inequalities instead of conic inequalities. For higher-dimensional conic constraints such as (5), differentiating (13) and (12) (instead of (9)) together would produce an invertible systems of equation to which the implicit function theorem can be applied. However, since deriving the derivatives involves complicated matrix differential calculus (Magnus and Neudecker, 2019), we omit its derivation in this paper.
A.2 Proof of Lemma 2
@@
A.3 Goal Generation for Local MPC
To generate goals that are considered more locally reachable yet far enough to be challenging for the planner, we utilize the following schemes. For the IiwaBimanual environment, we first prescribe some initial state $\bar{q}$ by sampling contact points on the surface of the object and solving inverse kinematics; then, the goals are sampled from the boundary of the action-only object motion set $\mathcal{M}^{\mathcal{A},\mathrm{o}}_{\mathbf{\Sigma},\kappa}(\bar{q},\bar{u})$ with $\bar{u}=\bar{q}^{\mathrm{a}}$. To make the goals sufficiently challenging, we use large ellipsoid radius for the purpose of sampling goals. The resulting goals have maximum $120^{\circ}$ rotation and $0.4m$ translation from the initial state (see Figure 11 for the distribution of goals). Similarly, we prescribe some grasp for AllegroHand and sample orientations that have about a $60^{\circ}$ difference from the initial state according to the axis-angle metric. With the above procedure, we sampled $1233$ pairs of initial states and goals for IiwaBimanual and $1000$ pairs for AllegroHand.
We summarize in Table 9 the average amount of translation and rotation between the initial and goal object configurations generated here. These would be the object pose tracking errors if the planning algorithm being evaluated did nothing.
@@ A.3 Goal Generation for Local MPC
Mean $\lVert p_{\text{goal}}-p_{0}\rVert$ [$m$] 152 16
Mean $\lvert\Delta\theta(Q_{\text{goal}},Q_{0})\rvert$ [mrad] 356 788
Table 9: Mean translation and rotation differences for pairs of initial and goal object states generated in this section.
A.4 Effect of Planning Horizon for Local MPC
We study how the planning horizon $T$ used in (36) affects the MPC errors. Specifically, for the IiwaBimanual system and the corresponding goals generated using the scheme in Section A.3, we run Algorithm 2 with planning horizon from $1$ to $5$, a maximum iteration of 22 ($T=1\dots 5$, $n_{\mathrm{max}}=2$ in Algorithm 1), and a MPC rollout horizon $H$ of $10$.
The results are summarized in Table 10. Surprisingly, we found that the planning horizon $T$ has little to no effect on the final error for MPC: our method is able to reach the goals within a tight tolerance for all values of $T$. We believe that this is due to two reasons: i) we have chosen the initial configurations and goals to be reasonably reachable using our trust region, and ii) the CQDC model allows large change in configurations within a single step, so we are able to get away with significantly less horizon compared to other MPC tasks such as dynamic locomotion.
@@ A.4 Effect of Planning Horizon for Local MPC
Table 10: Translation and rotation Errors for IiwaBimanual under different planning horizon $T$. The first row is the mean; the second row (numbers in brackets) is the standard deviation.
A.5 Hyperparameters for Local MPC
In Table 11, we summarize additional hyperparameters for local MPC. These parameters are used for all experiments involving Algorithm 2 and Algorithm 3, including the results in Figure 11 and Figure 12.
The MPC cost function is defined by the quadratic cost matrices $\mathbf{Q}$ and $\mathbf{R}$ in (37b). $\mathbf{Q}_{\mathrm{o}}$ is the submatrix of $\mathbf{Q}$ corresponding to the object. As we are primarily concerned with moving objects to goal poses, we set the robot-related entries in the cost matrix $\mathbf{Q}$ in (37) to 0, and rely on the action cost (37b) and the CTR constraint (37d) to stay close to the linearization points.
Other key hyperparameters include the CQDC smoothing parameter $\kappa$ and time step $h$. The choice of $\kappa$ and $h$ is a trade-off between the ease of planning and model accuracy. Larger $\kappa$ reduces the effect of contact dynamics smoothing: at $\kappa=10^{7}$, the smoothed CQDC dynamics (10) is very close to the non-smooth formulation (7). However, gradients of non-smooth dynamics are well-known to be less informative for local MPC. Similarly, smaller $h$ increases simulation fidelity, but also means MPC needs to take more steps to reach the goal. In practice, we tune these by sweeping across a range of values and selecting the best performers. Specifically, we tested $\kappa$ from $10^{2}$ and $10^{6}$ in orders or magnitude (e.g. $10^{2},10^{3},\dots$), and $h$ from $0.01$ to $0.1$.
To improve computational efficiency, we introduce a signed distance threshold, denoted as $\phi$-threshold in Table 11. When constructing a CTR or R-CTR, we only include feasibility constraints for contact pairs whose signed distance $\phi_{i}$ is less than this threshold. This technique effectively prunes contact pairs that are far from making contact, which reduces the number of constraints in the trust region and speeds up the MPC.
For our experiments, the values were chosen based on the object’s size: slightly smaller than the $0.28\mathrm{m}$ bucket diameter for IiwaBimanual and the $0.06\mathrm{m}$ cube edge length for AllegroHand. This is a very conservative choice. To speed up computation, we can compute a lower bound on the signed distance at contact $i$ at the next time step from $\phi_{i}$, $\mathbf{J}_{i}$ and $\mathbf{\Sigma}$. Excluding feasibility constraints corresponding to contact pairs whose lower bound is greater than 0 (or some small positive threshold) would significantly reduce problem size, thereby speeding up computation.
IiwaBimanual AllegroHand
$\mathbf{Q}_{\mathrm{o}}$ Rotation $0.1\mathbf{I}_{1}$ $\mathbf{I}_{3}$
Translation $\mathbf{I}_{2}$ $5\mathbf{I}_{3}$
$\mathbf{R}$ $01\mathbf{I}_{6} 0.0$$01\mathbf{I}_{16}$
$\kappa$ $10^{4}$ $10^{4}$
$h$ [s] $0.02$ $0.05$
$\phi$-threshold [m] 0.2 0.05
Table 11: Additional hyperparameters for MPC, used by both Algorithm 2 and Algorithm 3.
A.6 Hyperparameters for 2nd-order Dynamics Stabilization
We use the same $\mathbf{CtrTrajOpt}$ parameters from Table 2. Other important MPC hyperparameters are summarized in Table 12. For closed-loop IiwaBimanual experiments, we use larger $N$ for goals that are further away, while keeping $N\times H$ around $25$. In contrast, as slippage happens a lot more often on the AllegroHand system, we keep $N$ constant for all goals to give MPC more chances to recover.
To validate our choice of the rollout horizon $H$, we studied how different values of $H$ affect tracking errors in simulation, with results presented in Figure 22. We computed the Pearson correlation coefficient ($r$) between the and tracking error. The analysis revealed a negligible correlation for IiwaBimanual ($r=0.00082$) and a very weak one for AllegroHand (r=0.059r=-0.059). These findings confirm that performance is largely insensitive to $H$. Ultimately, our selection of $H$ is constrained by computational limitations. Because the robot must pause while the MPC replans, we selected values of $H$ that balanced the replanning frequency and duration of these pauses.
NN (Number of re-plans) $H$ (MPC rollout horizon)